# Towards Multi-modal Transformers in Federated Learning

### v1

@@ Abstract
Multi-modal transformers mark significant progress in different domains, but siloed high-quality data hinders their further improvement. To remedy this, federated learning (FL) has emerged as a promising privacy-preserving paradigm for training models without direct access to the raw data held by different clients. Despite its potential, a considerable research direction regarding the unpaired uni-modal clients and the transformer architecture in FL remains unexplored. To fill this gap, this paper explores a transfer multi-modal federated learning (MFL) scenario within the vision-language domain, where clients possess data of various modalities distributed across different datasets. We systematically evaluate the performance of existing methods when a transformer architecture is utilized and introduce a novel framework called Federated modality complementary and collaboration (FedCola) by addressing the in-modality and cross-modality gaps among clients. Through extensive experiments across various FL settings, FedCola demonstrates superior performance over previous approaches, offering new perspectives on future federated training of multi-modal transformers.
@@ 1 Introduction
Multi-modal transformers have led to remarkable advancements across a spectrum of downstream tasks [42, 2]. Nevertheless, training these models demands voluminous and high-quality data  [15, 30]. Although high-quality training datasets contain a wealth of information, acquiring such data is a resource-intensive task. Moreover, many of these datasets are often closely guarded by various entities and fall under diverse privacy regulations, resulting in data silos [21]. These silos pose remarkable challenges to further improvements in the field, as they prevent the leveraging of isolated datasets to enhance model training and performance.
In response to this challenge, federated learning (FL) [35] proposes a novel training paradigm typically designed to circumvent the need for direct access to raw data. FL enables a decentralized approach to model training. Specifically, a central server cyclically disseminates a global model to a selected pool of clients for local training. After the local training, the server receives the model updates from the clients, averages them with specific weights as the global aggregation, and applies the aggregated updates to the global model. The process continues until convergence. FL allows collaborative training on heterogenous data spread over geographically remote clients, while simultaneously preserving their data privacy, and offers a promising solution to the issue of data silos by facilitating the collaborative improvement of models.
@@
Current FL research predominantly addresses uni-modal scenarios and explores different aspects, such as local training [36, 27], personalization [48, 26, 3, 46, 34], aggregation [50, 19], and initialization [5, 38, 49]. Nonetheless, in domains with rich multi-modal data, such as healthcare and the Internet of Things [60, 53, 1], multi-modal federated learning (MFL) are gaining attention [4]. Studies have explored horizontal MFL in healthcare, which enables training clients that share identical input modalities (e.g., all clients have text and image) [53, 1, 6], and vertical MFL in the Internet of Things, which enables clients that possess a single modality from the same users, but it may differ between them (e.g., some clients have text-only data, while some have their corresponding image data) [60]. However, operating explicitly in either of the aforementioned settings excludes other uni-modal clients with unpaired data from participating in the MFL process. Unpaired uni-modal data has shown the capability to improve the multi-modal model under a centralized setting [2, 44]. Therefore, transferring knowledge from those uni-modal clients can further extend the scope of training data in FL, providing an unneglectable impact for MFL. But only a few studies have explored this transfer MFL setting [56, 32] shown in Fig. 1.
@@
To tackle these challenges, we propose a novel framework for the transfer MFL in both local training and global aggregation, leveraging the unified design of transformers across different modalities. During the local training, we propose a mixture strategy to leverage the modality-complementary information, along with a weight compression trick for efficient communication. During the global aggregation, we propose an aggregation with disaggregation strategy to aggregate general knowledge while maintaining specific knowledge for each local objective. With these collaboration techniques for both local training and global aggregation, we build and propose a novel framework, Federated modality Complementary and collaboration (FedCola), providing insights for FL on multi-modal transformers, and highlight the main contributions of this paper:
@@
    Cross-modality knowledge-sharing (§4.2). We illustrate that cross-modality knowledge-sharing can be achieved without directly accessing the data but only with the model weights, and provide new insights on convergence conditions of collaboration and quantified contributions.
@@ 2 Related Work
Federated Learning has emerged as a powerful approach for decentralized, privacy-preserving machine learning. In practice, heterogeneity in different aspects exists among clients, including statistical heterogeneity and system heterogeneity. The majority of current FL focuses on addressing the statistical heterogeneity for clients with data from the same domain, the same modality, but different distributions [36, 27, 22, 25]. Meanwhile, the other line of work focuses on enabling edge devices with limited resources to address the system heterogeneity [31, 17, 9]. Meanwhile, FL can also be categorized by the focus of the product, which can either be the performance of the global model on more generalized test data or the performance of the client models on their local data. These two categories lead to two separate directions of FL: global and personalized FL [34, 24, 46, 29, 11, 37]. In this paper, we explore and address the heterogeneity in the modality level to obtain a generalized multi-modal transformer, which falls into the scope of global FL.
@@
Vision Transformers in FL. Although vision transformers have become the de facto architecture in centralized computer vision tasks, their exploration under FL remains limited. Qu et al. [41] evaluate the performance of vision transformers under global FL settings and showcase their robustness against data heterogeneity. Meanwhile, Sun et al. [46] and Li et al. [24] study the personalization of vision transformers in personalized FL. Further, research with other aspects of vision transformers (e.g., pre-trained transformers) have also gained attention [47, 38, 5]. With a different perspective in this paper, we focus on the multi-modal capability and unified architecture of the transformer, paving the path for further large federated multi-modal transformers.
@@ 4.2 Complementary local training against cross-modality gap
However, if we directly send the updates of the modified local model, the upload communication cost will be almost doubled due to the involvement of more model parameters. To remedy this, we compute the equivalent weights for each layer after the local training as ${\bm{W}}_{\text{local}}+g{\bm{W}}_{\text{out}}$ as a compression trick since they are linear, so the total uploading size is reduced to the size of the original model. Regarding the download size, we further explore options to reduce the required model weights from the complementary modality in § 6.1. In this way, we provide the flexibility of the trade-off between performance and communication costs.
@@ 4.3 Collaborative aggregation against in-modality gap
Global aggregation in FL presents both opportunities and challenges. On one hand, aggregating model weights from diverse clients promises to enhance generalizability, while on the other, the inherent heterogeneity among clients poses significant obstacles to effective aggregation [35, 54]. While aggregation within the same modality constitutes a uni-aggregation problem, extensively studied within uni-modal FL, our focus shifts towards the collaboration dynamics among different modality/task combinations.
@@ 5 Experiments
Datasets. We follow the previous work’s setting to select the datasets for each modality type [56]. We use CIFAR-100 [23] as the image data for image clients and the AG NEWS [58] dataset as the text data for text clients. We evaluate the performance on two different multi-modal datasets, Flickr30k [39] and COCO Captions [8]. We sample $10,000$ images along with $50,000$ corresponding captions from Flickr30k noted as Flickr10k and $50,000$ image-text pairs from COCO Captions. To further evaluate the impact of the domain gap between uni-modal and multi-modal clients, we use two medical domain datasets to provide a larger domain gap. Specifically, we use OrganCMNIST [55] as the medical image dataset and the Medical Abstract [43] dataset as the medical text dataset. Classification tasks are performed on the uni-modal clients trained with cross-entropy loss, while cross-modal retrieval tasks are performed on multi-modal clients trained with contrastive loss [18].
@@
Table 1: Performance of (a) uni-modal methods, (b) multi-modal methods, and (c) our proposed method under different FL settings. The evaluation metric is the sum of top-1 recalls following previous settings.
@@ 5.2 Evaluation under different domain gaps
We investigate the performance of multi-modal methods compared to the FedAvg baseline relatively under these settings, and the result is shown in Fig. 4. We notice that FedCola consistently outperforms all other methods under all settings, demonstrating the robustness of our proposed method in handling various domain gaps. However, larger domain gaps do have a negative impact on the performance of FedCola, and we leave addressing the challenge of large domain gaps as future work. Similarly, FedIoT struggles the most in the setting with more gaps since its collaboration is based on the capability of multi-modal clients. When the domain gap is large, the multi-modal clients are not able to provide effective information to the uni-modal clients, leading to a performance drop. CreamFL, on the other hand, shows a more stable performance under different domain gaps since their collaboration is based on the public dataset, which is less affected by the domain gap, despite being unable to consistently outperform the uni-modal FedAvg baseline.
@@ 6.2 Fairness analysis
Unlike centralized training, where all data is owned by one data center aiming at one objective, the data in FL are from different clients for different local tasks. We further study the performance on uni-modal tasks and quantify the contributions of each type of client to the performance gain. Such an analysis provides insights into the fairness of the collaboration and an initial guideline for profit sharing among different clients [28, 45].
@@ 6.3 Scaling-up capability
To verify the scaling-up capability, we study the trend of the performance on multi-modal evaluation when more uni-modal datasets (i.e., CIFAR-100, AG NEWS, OrganCMNIST, and Medical Abstracts) are involved. As shown in Table 5, the performance increases when more generalized data is leveraged in FL, supporting the scaling-up capability of the proposed framework. More discussion is given in the supplementary material.
6.4 Visualization
@@ 7 Conclusion
In this work, we present FedCola, a novel framework tailored for transfer MFL focussing on transformers. By integrating complementary local training and collaborative aggregation techniques that solely depend on model parameters and dispense the need for public data, FedCola demonstrates improved performance over previous approaches under diverse FL scenarios, including different domain gaps and numbers of participating clients. Through extensive numerical experiments, our framework proves its effectiveness and offers new insights into the dynamics of collaboration, including conditions for convergence and the specific contributions of different client types. We envision that our contributions will encourage further investigation into multi-modal federated training paradigms, particularly to train large transformer architectures.
@@ 0.A.1 A general framework for transfer MFL
where $f_{ji}(w):=\frac{1}{D_{n}}\sum_{k=1}^{D}\mathbb{E}_{z_{k}\sim{\cal D}_{i}}{l(w;z_{k})}$ denotes the loss function evaluated on input datapoint at the $i^{\rm th}$ client, $M$ is the total number of modality types, and $N_{j}$ denotes the number of clients for the $j^{\rm th}$ modality. At the beginning of the training process, the server initializes global models, $w^{(0)}.$ The entire process runs for $T$ global communication rounds. At the beginning of each global round, $r$, each client, $i$, receives the global model, initializes it to $w_{i}^{0}=w^{(r)}$, and updates its local model in each iteration, $t$. We compactly write it:
@@ 0.A.2 Transfer MFL in our setting
We consider a transfer multi-modal FL setting in the vision-language domain with a total of $N$ clients. Based on the general framework, we describe our problem setup. Although (1) presents a general multimodal loss function, in our case, we are working with 3 modality types, so, $M=3$. Let $D_{\text{v}},D_{\text{l}},$ and $D_{\text{vl}}$ present the dataset for vision, text, and multimodal tasks, respectively. Note that, $D_{\text{vl}}=\{d:d=(v,l)\}$ contains vision and text pair as input, and define $D_{\text{vl}}(v):=P_{v}(D_{\text{vl}})$, $D_{\text{vl}}(l):=P_{l}(D_{\text{vl}}),$ where $P_{y}$ is the projection operator on the set $y.$ For training the vision models, we use $D_{\text{v}}\cup D_{\text{vl}}(v)$; for training the language models, we use $D_{\text{l}}\cup D_{\text{vl}}(l).$ At the beginning of the FL process, the server initializes global models, $w^{(0)}:={\left(w^{(0,\text{v})},w^{(0,\text{vl}_{\text{v}})},w^{(0,\text{vl}_{\text{l}})},w^{(0,\text{l})}\right)}^{\top}$ the set of parameters, and $\nabla w^{r}=\begin{pmatrix}\nabla w_{v}^{r}\;\nabla w_{l}^{r}\;\nabla w_{vl_{v}}^{r}\;\nabla w_{vl_{l}}^{r}\end{pmatrix}^{\top}.$
For local training, $\Omega_{\rm local}=\rm diag(\eta_{v}I_{|w_{v}|}\;\eta_{l}I_{|w_{v}|}\;\eta_{vl}I_{|w_{vl_{v}}|}\;\eta_{vl}I_{|w_{vl_{l}}|}),$ where $\eta_{v},\eta_{l},\eta_{vl}\geq 0$ are stepsizes for respective modality type. Note that, $|x|$ denotes the dimension of an arbitrary vector, $x,$ and $I_{|x|}$ is the identity matrix in the space $\mathbb{R}^{|x|\times|x|}.$ Let $n_{v},n_{l}$ and $n_{vl}$ clients be sampled uniformly at random from each $\{N_{j}\}_{j=1}^{3}$, respectively, in each training round. Hence, after $E$ local epochs, (0.A.2) is given as:

### v2

@@ Abstract
Multi-modal transformers mark significant progress in different domains, but privacy concerns on high-quality data hinder their further improvement. Federated learning (FL) has emerged as a promising privacy-preserving paradigm for training models without direct access to the raw data held by different clients. Despite its potential, a considerable research direction regarding the unpaired uni-modal clients and the transformer architecture in FL remains unexplored. To fill this gap, this paper explores a transfer multi-modal federated learning (MFL) scenario within the vision-language domain, where clients possess data of various modalities distributed across different datasets. We systematically evaluate the performance of existing methods when a transformer architecture is utilized and introduce a novel framework called Federated modality complementary and collaboration (FedCola) by addressing the in-modality and cross-modality gaps among clients. Through extensive experiments across various FL settings, FedCola demonstrates superior performance over previous approaches, offering new perspectives on future federated training of multi-modal transformers. Code is available at https://github.com/imguangyu/FedCola.
Keywords:
Federated Learning Multi-modal Learning Transformer
@@ 1 Introduction
Multi-modal transformers have led to remarkable advancements across a spectrum of downstream tasks [46, 2, 24]. Nevertheless, training these models demands voluminous and high-quality data  [16, 33]. Although high-quality training datasets contain a wealth of information, acquiring such data is resource-intensive. Moreover, many of these datasets are often closely guarded by various entities and fall under diverse privacy regulations, resulting in data silos [23]. These silos pose remarkable challenges to further improvements in the field, as they prevent the leveraging of isolated datasets to enhance model training and performance.
In response to this challenge, federated learning (FL) [38] proposes a novel training paradigm typically designed to circumvent the need for direct access to raw data. FL enables a decentralized approach to model training. Specifically, a central server cyclically disseminates a global model to a selected pool of clients for local training. After the local training, the server receives the model updates from the clients, averages them with specific weights as the global aggregation, and applies the aggregated updates to the global model. The process continues until convergence. FL allows collaborative training on heterogenous data spread over geographically remote clients while simultaneously preserving their data privacy and offers a promising solution to the issue of data silos by facilitating the collaborative improvement of models.
@@
Current FL research predominantly addresses uni-modal scenarios and explores different aspects, such as local training [40, 30], personalization [52, 29, 3, 50, 37], aggregation [54, 20], and initialization [5, 42, 53]. Nonetheless, in domains with rich multi-modal data, such as healthcare and the Internet of Things (IoTs) [64, 57, 1], multi-modal federated learning (MFL) are gaining attention [4]. Studies have explored horizontal MFL in healthcare, which enables training clients that share identical input modalities (e.g., all clients have text and image) [57, 1, 6], and vertical MFL in the IoTs, which enables clients that possess a single modality from the same users with multi-modal data (e.g., some clients may have text-only data, while some have their corresponding image data) [64]. However, operating explicitly in either of the settings excludes other uni-modal clients with unpaired data from participating in the MFL process. Unpaired uni-modal data has shown the capability to improve the multi-modal model under a centralized setting [2, 48]. Therefore, transferring knowledge from those uni-modal clients can further extend the scope of training data in FL, providing an important impact for MFL. But only a few studies have explored this transfer MFL setting [60, 35] shown in Fig. 1.
@@
To tackle these challenges, we propose a novel framework for the transfer MFL in both local training and global aggregation, leveraging the unified design of transformers across different modalities. During the local training, we propose a mixture strategy to leverage the modality-complementary information, together with a weight compression trick for efficient communication. During the global aggregation, we propose an aggregation with disaggregation strategy to aggregate general knowledge while maintaining specific knowledge for each local objective. With these collaboration techniques for both local training and global aggregation, we build and propose a novel framework, Federated modality Complementary and collaboration (FedCola), providing insights for FL on multi-modal transformers, and highlight the main contributions of this paper:
@@
    Cross-modality knowledge-sharing (§4.2). We illustrate that cross-modality knowledge-sharing can be achieved without directly accessing the data but only with the model weights and provide new insights on convergence conditions of collaboration and quantified contributions.
@@ 2 Related Work
Federated learning has emerged as a powerful approach for decentralized, privacy-preserving machine learning. In practice, heterogeneity in different aspects exists among clients, including statistical heterogeneity and system heterogeneity. Most current FL focuses on addressing the statistical heterogeneity for clients with data from the same domain, the same modality, but different distributions [40, 30, 25, 28, 66]. Meanwhile, the other line of work focuses on enabling edge devices with limited resources to address the system heterogeneity [34, 18, 9, 39, 21]. Meanwhile, FL can also be categorized by the focus of the product, which can either be the performance of the global model on more generalized test data or the performance of the client models on their local data. These two categories lead to two separate directions of FL: global and personalized FL [37, 27, 50, 32, 11, 41]. In this paper, we explore and address the heterogeneity in the modality level to obtain a generalized multi-modal transformer, which falls into the scope of global FL without system heterogeneity.
@@
Vision transformers in FL. Although vision transformers have become the de facto architecture in centralized computer vision tasks, their exploration under FL remains limited. Qu et al. [45] evaluate the performance of vision transformers under global FL settings and showcase their robustness against data heterogeneity. Meanwhile, Sun et al. [50] and Li et al. [27] study the personalization of vision transformers in personalized FL. Further, research with other aspects of vision transformers (e.g., pre-trained transformers) have also gained attention [51, 42, 5, 65]. With a different perspective in this paper, we focus on the multi-modal capability and unified architecture of the transformer, paving the path for further large federated multi-modal transformers.
@@ 4.2 Complementary local training against cross-modality gap
However, if we directly send the updates of the modified local model, the upload communication cost will be almost doubled due to the involvement of more parameters. To remedy this, we compute the equivalent weights for each layer after the local training as ${\bm{W}}_{\text{local}}+g{\bm{W}}_{\text{out}}$ as a compression trick since they are linear, so the total uploading size is reduced to the size of the original model. Regarding the download size, we further explore options to reduce the required model weights from the complementary modality in § 6.1. In this way, we provide the flexibility of the trade-off between performance and communication costs.
@@ 4.3 Collaborative aggregation against in-modality gap
Global aggregation in FL presents both opportunities and challenges. On one hand, aggregating model weights from diverse clients promises to enhance generalizability, while on the other, the inherent heterogeneity among clients poses significant obstacles to effective aggregation [38, 58, 14]. While aggregation within the same modality constitutes a uni-aggregation problem, extensively studied within uni-modal FL, our focus shifts towards the collaboration dynamics among different modality/task combinations.
@@ 5 Experiments
Datasets. We follow the previous work’s setting to select the datasets for each modality type [60]. We use CIFAR-100 [26] as the image data for image clients and the AG NEWS [62] dataset as the text data for text clients. We evaluate the performance on two different multi-modal datasets, Flickr30k [43] and COCO Captions [8]. We sample $10,000$ images along with $50,000$ corresponding captions from Flickr30k, noted as Flickr10k, and $50,000$ image-text pairs from COCO Captions. To further evaluate the impact of the domain gap between uni-modal and multi-modal clients, we use two medical domain datasets to provide a larger domain gap. Specifically, we use OrganCMNIST [59] as the medical image dataset and the Medical Abstract [47] dataset as the medical text dataset. Classification tasks are performed on the uni-modal clients trained with cross-entropy loss, while cross-modal retrieval tasks are performed on multi-modal clients trained with contrastive loss [19].
@@
Table 1: Performance of (a) uni-modal methods, (b) multi-modal methods, and (c) our proposed method under different FL settings. Arrows mean setting changes compared to Default. The evaluation metric is the sum of top-1 recalls following previous settings.
@@ 5.2 Evaluation under different domain gaps
We investigate the performance of multi-modal methods compared to the FedAvg baseline relatively under these settings, and the result is shown in Fig. 4. We notice that FedCola consistently outperforms all other methods under all settings, demonstrating the robustness of our proposed method in handling various domain gaps. However, larger domain gaps do have a negative impact on the performance of FedCola, and we leave addressing the challenge of large domain gaps as future work. Similarly, FedIoT struggles the most in the setting with more gaps since its collaboration is based on the capability of multi-modal clients. When the domain gap is large, the multi-modal clients are not able to provide effective information to the uni-modal clients, leading to a performance drop. CreamFL, on the other hand, shows a more stable performance under different domain gaps since their collaboration is based on the public dataset, which is less affected by the domain gap despite being unable to consistently outperform the uni-modal FedAvg baseline.
@@ 6.2 Fairness analysis
Unlike centralized training, where all data is owned by one data center aiming at one objective [14], the data in FL are from different clients for different local tasks. We further study the performance on uni-modal tasks and quantify the contributions of each type of client to the performance gain. Such an analysis provides insights into the fairness of the collaboration and an initial guideline for profit sharing among different clients [31, 49].
@@ 6.3 Scaling-up capability
To verify the scaling-up capability, we study the trend of the performance on multi-modal evaluation when more uni-modal datasets (i.e., CIFAR-100, AG NEWS, OrganCMNIST, and Medical Abstracts) are involved. As shown in Table 5, the performance increases when more generalized data is leveraged in FL, supporting the scaling-up capability of the proposed framework.
@@ 7 Conclusion
In this work, we present FedCola, a novel framework tailored for transfer MFL focussing on transformers. By integrating complementary local training and collaborative aggregation techniques that solely depend on model parameters and dispense the need for public data, FedCola demonstrates improved performance over previous approaches under diverse FL scenarios, including different domain gaps and numbers of participating clients. Our framework proves its effectiveness through extensive numerical experiments and offers new insights into the dynamics of collaboration, including conditions for convergence and the specific contributions of different client types. We envision that our contributions will encourage further investigation into multi-modal federated training paradigms, particularly to train large transformer architectures.
Acknowledgement
This work is partially supported by the NSF/Intel Partnership on MLWiNS under Grant No. 2003198.
Supplementary Material
The supplementary material is structured as follows:
  1. 1.
    In Section 0.A we sketch a general framework for transfer MFL (§0.A.1), then we adopt it to the transfer MFL setting in the vision-language (§0.A.2). We give the convergence guarantee in §0.A.3. Finally, in §0.A.4, we argue why multi-modal FL might work with diverse data modalities by providing a simple heuristic on the distribution of the multimodal data.
  2. 2.
    Section 0.B elaborates on more experimental details, including (i) implementation details, (ii) visualization of the data partitioning, (iii) stochasticity discussion, (iv) breakdown performance of the results in the main paper, (v) communication analysis, (vi) experiments under imbalanced total client numbers, and (vii) visualizations.
  3. 3.
    Section 0.C discusses the potential negative social impact and limitation of the proposed method.
@@ 0.A.1 A general framework for transfer MFL
where $f_{ji}(w)=\frac{1}{|D_{j}|}\sum_{i\in D_{j}}\mathbb{E}_{z\sim{\cal D}_{i}}{l(w;z)}$ denotes the loss function evaluated on input datapoint at the $i^{\rm th}$ client, with $D_{j}$ denoting the set of all clients whose data has non-empty $j$th modality, $M$ is the total number of modality types, and $N_{j}$ denotes the number of clients for the $j^{\rm th}$ modality. At the beginning of the training process, the server initializes global models, $w^{(0)}.$ The entire process runs for $T$ global communication rounds. At the beginning of each global round, $r$, each client, $i$, receives the global model, initializes it to $w_{i}^{0}=w^{(r)}$, and updates its local model in each iteration, $t$. We compactly write it:
@@ 0.A.2 Transfer MFL in our setting
We consider a transfer multi-modal FL setting in the vision-language domain with a total of $N$ clients. Based on the general framework, we describe our problem setup. Although (1) presents a general multimodal loss function, in our case, we are working with 3 modality types, so, $M=3$. Let $D_{\text{v}},D_{\text{l}},$ and $D_{\text{vl}}$ present the dataset for vision, text, and multimodal tasks, respectively. Note that, $D_{\text{vl}}=\{d:d=(v,l)\}$ contains vision and text pair as input, and define $D_{\text{vl}}(v):=P_{v}(D_{\text{vl}})$, $D_{\text{vl}}(l):=P_{l}(D_{\text{vl}}),$ where $P_{y}$ is the projection operator on the set $y.$ For training the vision models, we use $D_{\text{v}}\cup D_{\text{vl}}(v)$; for training the language models, we use $D_{\text{l}}\cup D_{\text{vl}}(l).$ At the beginning of the FL process, the server initializes global models, $w^{(0)}:={\left(w^{(0,\text{v})},w^{(0,\text{l})},w^{(0,\text{vl}_{\text{v}})},w^{(0,\text{vl}_{\text{l}})}\right)}^{\top}$ the set of parameters, and $\nabla w^{r}=\begin{pmatrix}\nabla w_{v}^{r}\;\nabla w_{l}^{r}\;\nabla w_{vl_{v}}^{r}\;\nabla w_{vl_{l}}^{r}\end{pmatrix}^{\top}.$
For local training, $\Omega_{\rm local}=\rm diag(\eta_{v}I_{|D_{v}|}\;\eta_{l}I_{|D_{l}|}\;\eta_{vl}I_{|D_{vl}(v)|}\;\eta_{vl}I_{|D_{vl}(l)|}),$ where $\eta_{v},\eta_{l},\eta_{vl}\geq 0$ are stepsizes for respective modality type. Note that, $|x|$ denotes the dimension of an arbitrary vector, $x,$ and $I_{|x|}$ is the identity matrix in the space $\mathbb{R}^{|x|\times|x|}.$ Let $n_{v},n_{l}$ and $n_{vl}$ clients be sampled uniformly at random from each $\{N_{j}\}_{j=1}^{3}$, respectively, in each training round. Hence, after $E$ local epochs, (3) is given as: